# A Stochastic Block Hypergraph model

### v1

@@ Abstract
We propose a simple model for a hypergraph generalization of the stochastic block model, using the clustering connection probability $P_{ij}$ between communities $i$ and $j$, and integrating explicitly the hyperedge formation process. Indeed, hyperedges are groups of nodes and we can expect that different real-world networks correspond to different formation processes of these groups and to different levels of homophily between nodes. We describe a simple model where we can explicitly introduce the hyperedge formation process and study its impact on the composition of hyperedges. We focus on the standard case where $P_{ij}=p\delta_{ij}+q(1-\delta_{ij})$, and when $0\leq q\leq p$, we show that the degree and hyperedge size distributions can be approximated by binomials with effective parameters that depend on the number of communities and on $q/p$. Also, the composition of hyperedges goes for $q=0$ from ‘pure’ hyperedges (comprising nodes belonging to the same community) to ‘mixed’ hyperedges that comprise nodes from different communities for $q=p$. We tested various formation processes and our results suggest that when they depend on the composition of the hyperedge, they tend to favor the dominant community and lead to hyperedges with a smaller diversity. In contrast, for formation processes that are independent from the hyperedge structure, we obtain hyperedges comprising a larger diversity of communities. The advantages of the model proposed here are its simplicity and flexibility that make it a good candidate for testing community-related problems, from their detection to their impact on various dynamics.
@@ I Introduction
The stochastic block model (SBM) is a simple model of a graph with communities. This generative model for data benefits from a ground truth for the communities, and has been widely used as a canonical model for community detection (see the review [1]). The standard community detection problem [2] is the statistical recovery problem of the clusters: from the graph, can we reconstruct the sets of nodes that describe the communities ? Depending on the parameter values there are different regimes such as impossible to recover, easy, etc. [3, 4, 5]). Most of these works concerned networks, but recent analysis of complex systems such as in systems biology [6], face-to-face systems [7], collaboration teams and networks [8, 9], ecosystems [10], the human brain [11, 12], document clusters in information networks, or multicast groups in communication networks [13], etc. showed that graphs provide a limited view, and in order to include higher-order interactions involving groups of more than two units, it is necessary to refer to hypergraphs [16, 14, 15]. These objects that are a natural extension of usual graphs, consists in allowing edges to connect an arbitrary number of nodes. These ‘hyperedges’ constructed over a set of vertices define what is called a hypergraph [17, 18]. More formally, a hypergraph is defined as $H=(V,E)$ where $V$ is a set of elements (the vertices or nodes) and $E$ is the set of hyperedges where each hyperedge is defined as a nonempty subset of $V$.
@@
Given the recent importance of these objects for describing cases that are not well described by graphs, it is natural to extend existing models. For instance, the random Erdos-Renyi graph can be extended in a number of ways such as in [19] (and references therein). Growing hypergraph models were also proposed in [20, 22, 21, 23] and in particular, introduced preferential attachment in this context [20]. In [23], random recursive hypergraphs grow by adding at each time step a vertex and a edge by joining the new vertex to a randomly chosen existing edge.
@@
Some other generalizations were proposed, such as a null model for hypergraphs with fixed node degree and edge dimensions [24, 21] which generalizes to hypergraphs the usual configuration model [25]. Other models for higher-order interactions were described in the review [16] such as bipartite models, exponential graph models, or motifs models, but it is fair to say that in general hypergraph modeling is less developed than its network counterpart. Most of these hypergraph models are introduced in the mathematical literature and are usually thought as immediate generalizations of classical graph models, where it is assumed that all hyperedges have the same size (and are called $k-$uniform hypergraphs when the size of hyperedges is $k$), which is a strong constraint.
@@
Concerning the SBM, some papers proposed an extension to hypergraphs (see for example [26, 27]). Most of the recent models such as in [28, 29, 30, 31], consider hypergraphs with fixed degree and/or edge sizes (which in general do not correspond to empirical observations). More precisely, in [26], the authors propose a generative approach to hypergraph clustering based on a degree-corrected hypergraph stochastic blockmodel. This model generates clustered hypergraphs with heterogeneous degree distributions and hyperedge sizes, based on a probability distribution over the space of possible hypergraphs. This is obviously a very general model, but difficult to use and to connect with real hypergraphs and processes at play in the formation and evolution of these structures. In [29], the authors study the problem of community detection in a random hypergraph model which they call the stochastic block model for $k-$uniform hypergraphs (a model introduced in [28]). Each hyperedge appears in the hypergraph independently with a probability depending on the community labels of the vertices involved in the hyperedge. Other recent works constructed random hypergraphs with communities [27, 32] and generalized the Poisson stochastic block model. In particular, in [27], the authors proposed a sampling approach where hyperedges are generated according to some probability and where nodes belong to different communities. As the authors of [32] noted, even if hypergraphs are widely adopted tools to examine systems with higher-order interactions, we are still lacking a theoretical understanding of their detectability limits in the community detection problem [30, 31, 33, 34, 35]. In this respect, it seems important to develop simple models of hypergraphs with communities in order to test various detection methods, but also to explore the impact of such a structure on various dynamics [36] such as diffusion [37], synchronization [38, 39] or spreading [40, 41, 42].
@@
Most of these models are however very general, difficult to code and are usually not well adapted to the case of growing hypergraphs. Indeed, these works take in general an ‘ensemble’ perspective where the hypergraph is an element chosen according to some distribution in a large set of hypergraphs. Here, we choose another setting, where we perform loops over the hyperedges and the nodes and ask if a given node joins with an existing hyperedge. The crucial ingredient here is thus the formation process of hyperedges. Hyperedges are in fact groups of nodes and we can envision that different real-world networks correspond to different formation processes of these groups. This is particularly relevant in the case of growing structures: a new node arrives in the system and will connect to existing hyperedges, or in other words, will join an existing group of nodes. We thus propose such a simple model where we can explicitly introduce the hyperedge formation process and study its impact on the hypergraph structure.
@@ II.2 A hypergraph generalization
The number $N=|V|$ of vertices is called the order of the hypergraph, and the number of hyperedges $M=|E|$ is usually called the size of the hypergraph. The size $m_{i}=|e_{i}|$ of an hyperedge $e_{i}$ is the number of its vertices, and the degree of a vertex is then simply given by the number of hyperedges to which it is connected. A simpler hypergraph considered in many studies is obtained when all hyperedges have the same cardinality $k$ and is called a $k$-uniform hypergraph (a $2$-uniform hypergraph is then a standard graph). A simple way to represent a hypergraph is by using its bipartite description where on one side we have nodes and on the other the hyperedges. A link in this bipartite representation describes then the membership of a node in a hyperedge (see for example Fig. 1)
@@
  • The minimal input is the number of nodes ($N$), the set of communities ($C_{1},C_{2}\ldots,C_{K}$), and the connection probability matrix $P_{ij}$ (here we will focus on the form given in Eq. 4).
@@
  • The code should be‘simple’ enough - in particular, it should explicitly display the formation process of hyperedges.
We now describe such a model. We start from $N$ vertices and a given number $E$ of hyperedges. We then perform a loop over the $N$ nodes and use the probability $\mathrm{Prob}(v\in e)$ to test if the node $v$ will be included in the hyperedge $e$. After a certain number of steps, a given hyperedge is of the form $e=\{u_{1},u_{2},\ldots,u_{n}\}$ where the community of $u_{l}$ is denoted by $C(u_{l})$. We now choose another node $v$ and we have to define the probability $\mathrm{Prob}(v\to e)$ that $v$ will belong to the hyperedge $e$. The main point here is that this framework allows us to specify the formation process of the hyperedge, or, in other words, what drives its composition.
There are several possible choices for this probability $\mathrm{Prob}(v\in e)$, as it could depend on the specific structure of the hyperedge $e$ and various distances between $v$ and $e$ (for example for spatial hypergraphs, the probability $\mathrm{Prob}(v\in e)$ could depend on the shortest or average distance from $v$ to $e$). At this stage we could also integrate some knowledge about group or team formation [51] and test their impact on the hypergraph structure. Here, we will focus on four simple models. For the first model ‘weighted’, we choose
$\mathrm{Prob}(v\to e)=\frac{1}{n}\sum_{j=1}^{n}P_{C(v)C(j)}$ (6)
@@
Here, we stop the construction of the hyperedge when all nodes are tested (in which case the probability matrix $P$ fixes the size of hyperedges), but other choices are possible. For example, we could stop the code when the size of all hyperedges reaches a value $k$ which is fixed (we then obtain what is called a $k-$uniform hypergraph), or when the size of the hyperedge reaches a random value $k$ distributed according to a distribution $P(k)$ (given empirically for example). In this paper, we focus on the simpler scenario of a fixed number of hyperedges that have a value fixed by the matrix $P$, which does not necessitate additional user-given constraints.
@@ III.1 Probability distribution for the degree and hyperedge sizes
In the case $q\neq p$, the probability that a node connects to an hyperedge will not be constant and will depend on the hyperedge structure. However, when $q<p$, we introduce a smaller probability to connect and we can expect a homogeneous behavior with probability $p$ but over a smaller effective subset of nodes of size $N^{*}$ and hyperedge sizes $E^{*}$ ($E^{*}$ for $P(k)$ and $N^{*}$ for $P(m)$). It is then reasonable to fit the degree and size distributions with binomials characterized by these effective parameters.
We show for different values of $q/p$, the degree distribution in Fig. 2, and the hyperedge size distribution in Fig. 3.
Figure 2: Degree distributions for $N=1000$, $E=200$, $p=\frac{30}{N}=0.03$, and $q/p=0,0.3,0.7,1$. The strategy used here is the weighted probability (Eq. 6). The distributions are fitted with binomial distributions of parameters $(p,N^{*})$.
Figure 3: Hyperedge size distributions for $N=1000$, $E=200$, $p=\frac{30}{N}=0.03$, and $q=0,0.3p,0.7p,p$. The strategy used here is the weighted probability (Eq. 6). The distributions are fitted with binomial distributions of parameters $(p,E^{*})$.
These figures show that even when $q\neq p$, these distributions can be approximated by binomials (which can however not be exact as the probability for a node to connect to a hyperedge depends on its composition and is in general not constant). We fit these distribution with binomials with effective parameters $N^{*}$ and $E^{*}$ and plot in Fig. 4 these parameters versus $q/p$.
Figure 4: Effective normalized parameters (top: $N^{*}/N$, and bottom: $E^{*}/E$) obtained by fitting with binomials the degree and size distributions. Here the hypergraphs are obtained for $N=1000$, $E=200$, $p=\frac{30}{N}=0.03$ and the weighted model. The lines represent linear fits of the form $a(q/p)+b$ where $a=0.73$ and $b=0.27$ for both plots, in agreement with our argument (Eq. 12.
@@
These results show that for this model, the degree and size distributions are therefore not very complex and can reasonably be approximated by binomials.
@@ III.2 Composition of hyperedges
In order to characterize the composition of a hyperedge, we denote by $n_{i}$ the number of nodes from community $i$ (where $i\in[[1,K]]$, and a natural quantity that describes this distribution is the Gini coefficient defined as [50]
@@
The influence of $q$ and $p$ on the composition of the hyperedges is illustrated on Fig. 5, where we see that for low values of $q/p$ , the hyperedges tend to be pure, which reveal the community structure of the hypergraph. Interestingly, we note here the existence of hyperedges at the interface of the different communities and that are characterized by a smaller value of their Gini coefficient. For larger values, such as $q/p=0.5$ (c), hyperedges become even more mixed, that tend to make the detection of the underlying community more complicated, if not impossible. When $p=q$ (d), the Gini coefficient $G$ is close to 0 on average, as illustrated by the darkness of the hyperedges. In other words, most hyperedges are mixed, as the processus of formation of the hyperdgees is community agnostic for this value.
@@
Figure 5: Drawing of four generated hypergraphs with $N=E=80$, $K=4$, $p=0.1$, for the $majority$ strategy, and for (a) $q=0$, (b) $q=0.01$, (c) $q=0.05$, and (d) $q=0.1$. Each community is of the same size with $N/K=20$ nodes. Hyperedges are drawn using a bipartite representation such as in Figure 1 using a force-directed layout implemented in D3 [49], with circles encoding nodes and squares hyperedges. The size of nodes and hyperedges encode their respective degree and size. The color of nodes encodes their community (each color is one community), while the color of hyperedges represents their Gini values.
@@
Figure 6: Distribution of $G$ for various values of $q$ ranging from 0 to $p$, with $p=0.1$, $N=2000$, $E=200$, $K=4$, and for the different strategies.
When $q/p$ decreases, the probability to connect together nodes from different communities decreases and most hyperedges will be pure (or close to pure) which corresponds to $P(G)\sim\delta(G-1)$ (where $\delta$ is the delta distribution). We note that the convergence to $\delta$ is not the same for all strategies: for the ’majority’ strategy, the progression of $G$ seems to follow a steady pattern from low to high values when increasing $q/p$, while the distribution of $G$ has a high variation rapidly for the other strategies. For $q/p$ increasing, hyperedges become more and more mixed until $q/p\sim 1$ and all communities are equally represented leading to $G\approx 0$.
@@
Figure 7: $\overline{G}$ (average of $G$) for $q$ ranging from 00 to $p$, with $E=200$ and different $N$ and for the strategy ‘weighted’.
@@
Figure 8: Average Gini coefficient $\overline{G}(q/p=1)$ versus $\frac{1}{\sqrt{\frac{N}{E}}}$ with $E=200$, and $p=q=0.03$.
@@
Figure 9: $\overline{G}$ (average of $G$) for $q$ ranging from 00 to $p$, with $E=200$ and $N=1,000$ for different strategies.
@@
Figure 10: Normalized standard deviation $\delta$ of $G$ for $q$ ranging from 00 to $p$ and for different $N$. From left to the right, we show the result for the different strategies ($p=0.03$, $E=200$)..
We observe on this figure that $\delta$ is increasing with $q/p$. For $q\approx 0$, all hyperedges are almost pure and when $q$ is increasing, fluctuations increase in general quickly. For $q=p$, all strategies lead to the same fluctuation. We observe here two remarkable facts. First, for all strategies, except for the ‘max’ one, $\delta$ is a monotonously increasing function of $q/p$. Indeed, for the ‘max’ strategy, there is an optimum here around $q/p\approx 0.2$. At this point, the heterogeneity among hyperedges is the largest. This example shows that the largest heterogeneity is not necessarily obtained for $q=p$ but depends strongly on the hyperedge formation process.

### v2

@@ Abstract
The stochastic block model is widely used to generate graphs with a community structure, but no simple alternative currently exists for hypergraphs, in which more than two nodes can be connected together through a hyperedge. We discuss here such a hypergraph generalization, based on the clustering connection probability $P_{ij}$ between nodes of communities $i$ and $j$, and that uses an explicit and modulable hyperedge formation process. We focus on the standard case where $P_{ij}=p\delta_{ij}+q(1-\delta_{ij})$ when $0\leq q\leq p$ ($\delta_{ij}$ is the Kronecker symbol). We propose a simple model that satisfies three criteria: it should be as simple as possible, when $p=q$ the model should be equivalent to the standard hypergraph random model, and it should use an explicit and modulable hyperedge formation process so that the model is intuitive and can easily express different real-world formation processes. We first show that for such a model the degree distribution and hyperedge size distribution can be approximated by binomial distributions with effective parameters that depend on the number of communities and $q/p$. Also, the composition of hyperedges goes for $q=0$ from ‘pure’ hyperedges (comprising nodes belonging to the same community) to ‘mixed’ hyperedges that comprise nodes from different communities for $q=p$. We test various formation processes and our results suggest that when they depend on the composition of the hyperedge, they tend to favor the dominant community and lead to hyperedges with a smaller diversity. In contrast, for formation processes that are independent from the hyperedge structure, we obtain hyperedges comprising a larger diversity of communities. The advantages of the model proposed here are its simplicity and flexibility that make it a good candidate for testing community-related problems, such as their detection, impact on various dynamics, and visualization.
@@ I Introduction
The stochastic block model (SBM) is a simple model of a graph with communities. This generative model for data benefits from a ground truth for the communities, and has been widely used as a canonical model for community detection (see the review [1]). The standard community detection problem [2] is the statistical recovery problem of the clusters: from the graph, can we reconstruct the sets of nodes that describe the communities? Depending on the parameter values there are different regimes such as impossible to recover, easy, etc. [3, 4, 5]. Most of these works concerned networks, but recent analysis of complex systems such as in systems biology [6], face-to-face systems [7], collaboration teams and networks [8, 9], ecosystems [10], the human brain [11, 12], document clusters in information networks, or multicast groups in communication networks [13], etc. showed that graphs provide a limited view, and in order to include higher-order interactions involving groups of more than two units, it is necessary to refer to hypergraphs [17, 15, 16]. These objects that are a natural extension of usual graphs, consist in allowing edges to connect an arbitrary number of nodes. These ‘hyperedges’ constructed over a set of vertices define what is called a hypergraph [14, 18]. More formally, a hypergraph is defined as $H=(V,E)$ where $V$ is a set of elements (the vertices or nodes) and $E$ is the set of hyperedges where each hyperedge is defined as a nonempty subset of $V$.
@@
Given the recent importance of these objects for describing cases that are not well described by graphs, it is natural to extend existing models. For instance, the random Erdos-Renyi graph can be extended in a number of ways, such as in [19] (and references therein). Growing hypergraph models were also proposed in [20, 23, 22, 24, 21] and in particular, introduced preferential attachment in this context [20, 21]. In [24], random recursive hypergraphs grow by adding at each time step a vertex and an edge by joining the new vertex to a randomly chosen existing edge.
@@
Some other generalizations were proposed, such as a null model for hypergraphs with fixed node degree and edge dimensions [25, 22] which generalizes to hypergraphs the usual configuration model [26]. Other models for higher-order interactions were described in the review [17] such as bipartite models, exponential graph models, or motifs models, but it is fair to say that in general hypergraph modeling is less developed than its network counterpart. Most of these hypergraph models are introduced in the mathematical literature and are usually thought of as immediate generalizations of classical graph models, where it is assumed that all hyperedges have the same size (and are called $k-$uniform hypergraphs when the size of hyperedges is $k$), which is a strong constraint.
@@
Concerning the SBM, some papers proposed an extension to hypergraphs (see for example [27, 28]). Most of the recent models such as in [29, 31, 32, 33], consider hypergraphs with fixed degree and/or edge sizes (which in general do not correspond to empirical observations). More precisely, in [27], the authors propose a generative approach to hypergraph clustering based on a degree-corrected hypergraph stochastic blockmodel. This model generates clustered hypergraphs with heterogeneous degree distributions and hyperedge sizes, based on a probability distribution over the space of possible hypergraphs. This is obviously a very general model, but difficult to use and to connect with real hypergraphs and processes at play in the formation and evolution of these structures. In [31], the authors study the problem of community detection in a random hypergraph model which they call the stochastic block model for $k-$uniform hypergraphs (a model introduced in [29]). Each hyperedge appears in the hypergraph independently with a probability depending on the community labels of the vertices involved in the hyperedge. Other recent works constructed random hypergraphs models with communities and generalized the LFR community detection benchmark[30] or the Poisson stochastic block model[28, 34]. In particular, in [28], the authors proposed a sampling approach where hyperedges are generated according to some probability and where nodes belong to different communities. As the authors of [34] noted, even if hypergraphs are widely adopted tools to examine systems with higher-order interactions, we are still lacking a theoretical understanding of their detectability limits in the community detection problem [32, 33, 53, 38, 39]. In this respect, it seems important to develop simple models of hypergraphs with communities in order to test various detection methods, but also to explore the impact of such a structure on various dynamics [40] such as diffusion [41], synchronization [42, 43] or spreading [44, 45, 46].
@@
Most of these models are however very general, difficult to code, and are usually not well adapted to the case of growing hypergraphs. Indeed, these works take, in general an ‘ensemble’ perspective where the hypergraph is an element chosen according to some distribution in a large set of hypergraphs. Here, we choose another setting, where we perform loops over the hyperedges and the nodes and ask if a given node joins with an existing hyperedge. The crucial ingredient here is thus the formation process of hyperedges. Hyperedges are, in fact, groups of nodes and we can envision that different real-world networks correspond to different formation processes of these groups. This is particularly relevant in the case of growing structures: a new node arrives in the system and will connect to existing hyperedges, or in other words, will join an existing group of nodes. We thus propose such a simple model where we can explicitly introduce the hyperedge formation process and study its impact on the hypergraph structure.
@@ II.2 A hypergraph generalization
The number $N=|V|$ of vertices is called the order of the hypergraph, and the number of hyperedges $M=|E|$ is usually called the size of the hypergraph. The size $m_{i}=|e_{i}|$ of a hyperedge $e_{i}$ is the number of its vertices, and the degree of a vertex is then simply given by the number of hyperedges to which it is connected. A simpler hypergraph considered in many studies is obtained when all hyperedges have the same cardinality $k$ and is called a $k$-uniform hypergraph (a $2$-uniform hypergraph is then a standard graph). A simple way to represent a hypergraph is by using its bipartite description where on one side we have nodes and on the other the hyperedges. A link in this bipartite representation describes then the membership of a node in a hyperedge (see for example Fig. 1)
@@
  • The model should be as simple as possible, where the minimal input is the number of nodes ($N$), the number of hyperedges ($E$), the set of communities ($C_{1},C_{2}\ldots,C_{K}$), and the connection probability matrix $P_{ij}$ (here we will focus on the form given in Eq. 1).
@@
  • The hyperedge formation process should be explicit and modulable, to express different types of hyperedge formation scenarios.
We now describe such a model. We start from $N$ vertices and a given number $E$ of hyperedges. For each hyperedge $e$, we perform a loop over the $N$ nodes and use the probability $\mathrm{Prob}(v\in e)$ to test if the node $v$ will be included in the hyperedge $e$. After a certain number of steps, a given hyperedge is of the form $e=\{v_{1},v_{2},\ldots,v_{n}\}$ where the community of $v_{l}$ is denoted by $C(v_{l})$. When adding another node $v$, we have to define the probability $\mathrm{Prob}(v\to e)$ that $v$ will belong to the hyperedge $e$. The main point here is that this framework allows us to specify the formation process of the hyperedge, or, in other words, what drives its composition.
There are several possible choices for this probability $\mathrm{Prob}(v\in e)$, as it could depend on the specific structure of the hyperedge $e$ and various distances between $v$ and $e$ (for example for spatial hypergraphs, the probability $\mathrm{Prob}(v\in e)$ could depend on the shortest or average distance from $v$ to $e$). At this stage, we could also integrate some knowledge about group or team formation [56] and test their impact on the hypergraph structure. Here, we will focus on four simple models. For the first model ‘weighted’, we choose
$\mathrm{Prob}(v\to e)=\frac{1}{|e|}\sum_{u\in e}^{n}P_{C(u)C(v)}$ (3)
@@
These strategies could reflect different real-world interaction processes. For example, a person (node) joining a discussion group (hyperedge) may reflect the ‘weighted’ or the ‘frequent’ strategies, where the person gauges the frequency of people he knows in the group before joining it or not. In contrast, in scenarios where nodes have some notion of risk by joining hyperedges, the ‘min’ strategy could reflect better the behavior of the agents.
@@
Here, we stop the construction of the hyperedge when all nodes are tested (in which case the probability matrix $P$ fixes the size of hyperedges), but other choices are possible. For example, we could stop the code when the size of all hyperedges reaches a value $k$ which is fixed (we then obtain what is called a $k-$uniform hypergraph), or when the size of the hyperedge reaches a random value $k$ distributed according to a distribution $P(k)$ (given empirically for example). In this paper, we focus on the simpler scenario of a fixed number of hyperedges that have a value fixed by the matrix $P$, which does not necessitate additional user-given constraints. In all cases, since we iterate over the nodes and hyperedges in a nested loop, the algorithm’s complexity is of $O(NE)$. When $N\sim E$, this can be approximated to $O(N^{2})$. We ran a small benchmark shown in Figure 2 that confirms the quadratic complexity of the algorithm when $N\sim E$.
Figure 2: Benchmark of the time (in seconds) needed to run the algorithm (a) for different values of N, with $N=E$, $K=4$, $p=\frac{100}{N}$, $q=0.4p$, using the $weighted$ strategy (each community is of the same size), and a quadratic fit (b) that shows the quadratic complexity of the algorithm ($f(N)=aN^{2}+bN+c$ with $a=1.0110^{-6}$, $b=-9.810^{-4}$, $c=0.34$).
@@ III.1 Probability distribution for the degree and hyperedge sizes
In the case $q\neq p$, the probability that a node connects to a hyperedge will not be constant and will depend on the hyperedge structure. When $q<p$, we introduce a smaller probability to connect, and we can expect a homogeneous behavior with probability $p$ but over a smaller effective subset of nodes of size $N^{*}$ and hyperedge sizes $E^{*}$ ($E^{*}$ for $P(k)$ and $N^{*}$ for $P(m)$). It is then reasonable to fit the degree and size distributions with binomials characterized by these effective parameters.
We show for different values of $q/p$, the degree distribution in Fig. 3 (a), and the hyperedge size distribution in Fig. 3 (b).
Figure 3: Degree (a) and hyperedge size (b) distributions computed from $100$ hypergraphs with $N=1000$, $E=200$, $K=4$, $p=\frac{30}{N}=0.03$, and $q/p=0,0.3,0.7,1$. The strategy used is the weighted probability (Eq. 3). The distributions are fitted with binomial distributions of parameters $(p,N^{*})$.
This shows that even when $q\neq p$, these distributions can be approximated by binomials (which can, however, not be exact as the probability for a node to connect to a hyperedge depends on its composition and is generally not constant).
We fit these distributions with binomials with effective parameters $N^{*}$ and $E^{*}$ and plot in Fig. 4 these parameters versus $q/p$.
Figure 4: Effective normalized parameters (top: $N^{*}/N$, and bottom: $E^{*}/E$) obtained by fitting with binomials the degree and size distributions. The distributions are computed using $100$ hypergraphs generated with $N=1000$, $E=200$, $K=4$, $p=\frac{30}{N}=0.03$ and the weighted formation process. The lines represent linear fits of the form $a(q/p)+b$ where $a=0.73$ and $b=0.27$ for both plots, in agreement with our argument (Eq. 9).
@@
These results show that for this model, the degree and size distributions are therefore not very complex and can reasonably be approximated by binomials. In our simulations, the binomial fit indeed works well for all strategies and binomial distribution were indeed observed in real world networks such as in [53]. Other studies (see for example [36, 37] and references therein) show empirical results with broad distributions for example. In order to reproduce this type of distributions, we certainly have to bring a crucial modification to the model (such as another strategy for example), but we note that the hyperedge distribution drifts away from a binomial distribution for the $min$ and $max$ strategies when $q\ll p$. In this case, the hyperedge distribution starts to take a bimodal shape, illustrated in Figure 5.
Figure 5: Hyperedge size distribution computed from $100$ hypergraphs with $N=1000$, $E=200$, $K=4$ $p=0.1$, and $q/p=0.1,0.01$, with the $min$ strategy.
This can be explained by the following ‘threshold’ mechanism: with the $min$ strategy, most hyperedges will be pure, as adding a node from another community is associated with the very low probability $q$. However, when this happens, the hyperedge becomes mixed, and from this moment, every node tested from any community will have a very low probability of joining the hyperedge (caused by the very low $q$). We, therefore, see very different sizes for pure and mixed hyperedges. We see a similar threshold mechanism for the $max$ strategy, but the other way around: mixed hyperedges have a higher size on average than pure hyperedges.
@@ III.2 Composition of hyperedges
To characterize the composition of a hyperedge, we denote by $n_{i}$ the number of nodes from community $i$ (where $i=1,2,\dots,K$) in the hyperedge, and a natural quantity that describes this distribution is the Gini coefficient defined as [55]
@@
The influence of $q$ and $p$ on the composition of the hyperedges is illustrated on Fig. 6, where we see that for low values of $q/p$ , the hyperedges tend to be pure, which reveal the community structure of the hypergraph. Interestingly, we note here the existence of hyperedges at the interface of the different communities and that are characterized by a smaller value of their Gini coefficient. For larger values, such as $q/p=0.5$ (c), hyperedges become even more mixed, that tend to make the detection of the underlying community more complicated, if not impossible. When $p=q$ (d), the Gini coefficient $G$ is close to 0 on average, as illustrated by the darkness of the hyperedges. In other words, most hyperedges are mixed, as the process of formation of the hyperedges is community agnostic for this value.
@@
Figure 6: Drawing of four generated hypergraphs with $N=E=80$, $K=4$, $p=0.1$, for the $majority$ strategy, and for (a) $q=0$, (b) $q=0.01$, (c) $q=0.05$, and (d) $q=0.1$. Each community is of the same size with $N/K=20$ nodes. Hyperedges are drawn using a bipartite representation such as in Figure 1 using a force-directed layout implemented in D3 [54], with circles encoding nodes and squares hyperedges. The size of nodes and hyperedges encode their respective degree and size. The color of nodes encodes their community (each color is one community), while the color of hyperedges represents their normalized Gini ($G$) values.
@@
Figure 7: Distribution of $G$ for various values of $q$ ranging from 0 to $p$, with $p=0.1$, $N=2000$, $E=200$, $K=4$, and for the different formation strategies. $100$ hyperedges are generated for each combination of parameters.
When $q/p$ decreases, the probability to connect together nodes from different communities decreases and most hyperedges will be pure (or close to pure) which corresponds to $P(G)\sim\delta(G-1)$ (where $\delta$ is the Dirac delta distribution). We note that the convergence to $\delta$ is not the same for all strategies: for the ’majority’ strategy, the progression of $G$ seems to follow a steady pattern from low to high values when increasing $q/p$, while the distribution of $G$ has a high variation rapidly for the other strategies. For $q/p$ increasing, hyperedges become more and more mixed until $q/p\sim 1$ and all communities are equally represented leading to $G\approx 0$.
We also test how the node traversal order may affect the Gini distribution, and, therefore, the structure of the generated hypergraphs. We review three cases: a) the node ordering is random for each hyperedge, b) the node ordering is random but the same for each hyperedge (‘fixed’), and c) the nodes are ordered given their community. The Gini distribution is shown in Figure 8 for the different ordering and hyperedge formation strategies. We can see that the community ordering slightly shifts the distribution to higher values, i.e., hyperedges are overall purer, for the min formation process. This is predictable, as once all nodes of community ‘A’ are tested for a hyperedge $e$, the nodes of the following communities will have a lower probability of entering the hyperedge since several nodes of community ‘A’ will already be in $e$. In practice, this node traversal order does not make much sense in a modeling scenario, and we advise using the random strategy for most cases.
Figure 8: Distribution of $G$ for different node traversal strategies and for different hyperedge formation processes, with $p=0.03$, $q=0.4p$ $N=1000$, $E=200$, $K=4$.
@@
Figure 9: Average $\overline{G}$ computed over $100$ hypergraphs for $q$ ranging from 00 to $p$, with $E=200$, different values of $N$, and for the ‘weighted’ strategy formation process.
@@
Figure 10: Average Gini coefficient $\overline{G}(q/p=1)$ versus $1/\sqrt{N/E}$ with $E=200$, and $p=q=0.03$. Each $\overline{G}$ is computed over $100$ generated hypergraphs.
@@
Figure 11: $\overline{G}$ (average of $G$) computed on 100 hypergraphs for $q$ ranging from 00 to $p$, with $E=200$ and $N=1,000$ for different strategies.
@@
Figure 12: Normalized standard deviation $\Delta$ of $G$ for $q$ ranging from 00 to $p$ and for different $N$. From left to right, we show the result for the different hyperedge formation strategies ($p=0.03$, $E=200$, $100$ generated hypergraphs for each set of parameters).
We observe in this figure that $\Delta$ is increasing with $q/p$. For $q\approx 0$, all hyperedges are almost pure and when $q$ is increasing, fluctuations increase in general quickly. For $q=p$, all strategies lead to the same fluctuation. We observe here two remarkable facts. First, for all strategies, except for the ‘max’ one, $\Delta$ is a monotonously increasing function of $q/p$. Indeed, for the ‘max’ strategy, there is an optimum here around $q/p\approx 0.2$. At this point, the diversity in hyperedge composition is the largest: there are a significant number of hyperedges tending toward a pure composition as $q$ is low, but hyperedges are mixing themselves more rapidly than with other strategies, thus creating a state where pure and mixed hyperedges coexist more than other strategies, explaining the optimum of $\Delta$. This example shows that the largest heterogeneity is not necessarily obtained for $q=p$ but depends strongly on the hyperedge formation process.